In a Mathematics contest, a contestant was awarded 6 points for each correct answer but had 4 points deducted for each incorrect answer.After answering 20 questions, Meina earned 40 points. How many questions did she answer correctly?
(algebra method / Guess and check is not allow)

Answer by nerdybill(7384)

Let x = number of correct of answers
then
20-x = number of incorrect answers
.
6x - 4(20-x) = 40
6x - 80 + 4x = 40
10x - 80 = 40
10x = 120
x = 12